OpenStudy (anonymous):

What is the value of 125^-1/3?

OpenStudy (anonymous):

cube root of 1/125

OpenStudy (anonymous):

\[ \frac{1}{125^{1/3}} = \frac{1}{5}\]

OpenStudy (anonymous):

1/5